Obtain an expression of probability for finding the free particle at position x confined in a box of length a
1
Expert's answer
2020-04-21T19:12:00-0400

The wave-function of a free particle at position x confined in 1-dimensional a box of length a can be found from Schrodinger equation:


"\psi=\sqrt{\frac{2}{a}}\text{ sin}\bigg[\frac{\pi n}{a}x\bigg]"